# SoK: Consensus for Fair Message Ordering

### v1

@@ IV-C Different Processing Approaches for Leader-based and Leaderless Consensus
Figure 5: Fair consensus factory processing leader-based consensus (top) and leaderless consensus (bottom) respectively
In Leader-based consensus, only the leader proposes an ordered set of messages and communicates to the peers, therefore, the factory simply removes the propose phase and replaces it with local rule executions. However, in a leaderless consensus, all nodes can propose in a round. These proposals are aggregated to derive a final ordered message set to commit. Therefore, when handling leaderless consensus, the factory removes both the propose phase and the aggregation phase, turning them into the same synchronize-and-execute flow as in the leader-based consensus.
In both leader-based and leaderless consensus, state synchronization can be integrated into the communication during leader election or in the message exchanges that signify the start of a new consensus round; in leader-based consensus, it can also be integrated into the leader election phase. In both cases, state synchronization does not introduce a new round of communication, which has little impact on the latency.
IV-D Limitation: Local Reordering for Messages in Less Than f+1 Local Message Pools
In practice, message senders of distributed ledgers usually send messages to only one node in the network (Figure 6 left). In this case, or more generally if a message only exists in less than $f+1$ nodes, then it is impossible to guarantee the message gets synchronized in the synchronization stage, Because the message can be held by the Byzantine nodes to not share with their peers, or they can collectively make a false claim of the message order.
A proposed workaround before the synchronization stage is to ask external message senders to communicate messages to all parties [23, 1] (Figure 6 right). This is in fact asking for an external synchrony assumption [20] such that messages from external senders are delivered to all nodes in a known time bound, which makes it impractical.
Figure 6: Two message dissemination modes
A practical existing solution for local reordering is to apply the blind ordering restriction. Although Byzantine nodes can theocratically still locally reorder messages, they are not motivated to do so if they are not able to censor the messages and derive a profitable strategy to reorder.
V Case Study: Themis
@@
Themis [1] is a FIFO via voting ordering protocol that is recognized as a potential solution to the blockchain order-fairness problem [14, 12, 18, 79]. In particular, Chen et al. [79] present performance optimization for the execution ordering algorithm used in Themis. We pick up this state-of-the-art consensus and apply the fair consensus factory framework to demonstrate the framework easily understands a fair ordering protocol and potentially provides performance optimizations.
V-A Understanding the Design of Themis
The core ordering algorithm for Themis is FIFO via voting algorithm that achieves batch order fairness. Given a set of unordered messages and the metadata of local message ordering from $t$ nodes, it orders the input messages by constructing a directed graph that contains information of, first, which messages should be included in the output, second, the order of messages. Compared to a previous similar proposal [20], Themis allows the ordering algorithm to output messages without seeing the complete batch (i.e. a complete Condorcet cycle), which reduces latency and provides stronger liveness.
To execute the ordering rule in a consensus, Themis uses Hotstuff [80] as the base consensus protocol and extends it to add a fair ordering feature. Figure 7 describes the Themis process in a round.
Figure 7: Illustration of the propose and pre-commit phase of Themis process. The system consists of five nodes $n_{1},...,n_{5}$, out of which $n_{1}$ is the leader of this round. Propose, pre-commit, commit, and decide phases are illustrated in the Hotstuff consensus [80]. Highlighted are the additional operations for fair ordering.
Although Themis plugs in a synchronization stage to Hotstuff, it only asks nodes to provide local states to the leader at the first round of communication. Note that nodes still need simultaneously to send the leader a New-View message as in the original Hotstuff, signifying the start of a new consensus round. After that, the leader is able to execute the ordering algorithm and derive the ordered message set for this round. But to convince the peers to agree on this ordered message set, the leader has to provide all local states to the peers so that they can run the algorithm in the subsequent step to verify the execution result of the leader.
V-B Optimizing Themis with Fair Consensus Factory
If the fair consensus factory is used to redesign Themis using its proposed ordering algorithm and Hotstuff as the base consensus, the result is as Figure 8.
Figure 8: Illustration of optimized propose and pre-commit phase Themis process. The propose phase is compressed into an all-to-all message exchange instead of a two-stage communication.
This design reduces the two-stage synchronization (all-to-leader, leader-to-all) in the original Themis protocol into a single-stage all-to-all synchronization. The latency optimization can be proven with existing latency comparisons between the two consensus protocols Hotstuff and PBFT [81]. Similar to the difference between Themis and the optimized Themis we propose, a key difference between Hotstuff and PBFT is that Hotstuff employs two-stage communication to reach agreements while PBFT employs single-stage all-to-all communication. The single-stage all-to-all communication mode has shown to have a lower latency [81]. Moreover, allowing all nodes to locally execute rules at the same stage eliminates the latency introduced by the separated leader and peer rule execution stages.
VI Fair Ordering Research Roadmap
@@
VI-A How to Choose Fair Ordering Rules
With these different notions of fairness, there are few discussions on which one is the “better” ordering rule, or which one is more suitable for distributed ledgers powering which applications. A possibility is that the sequencing rules should be application-specific [82], e.g., first-come-first-serve is more suitable for financial applications to prevent MEV and reduce latency [4]. However, outside of financial applications, it is argued that the current most popular fee-based prioritization approach [83] results in higher overall social welfare [84] compared to FIFO. Also, there are few discussions or studies on whether random ordering is indeed fair and whether there are scenarios that suit random ordering.
VI-B Public Verifiability of Order Fairness
In existing fair ordering proposals, external parties are unable to verify whether the ordering results are fair. Without such verifiable fairness, there is a risk for the majority of nodes in the system to collude and corrupt only the order fairness in the consensus, which is non-detectable without verifiability. That is, from the view of the message senders, it is indistinguishable whether a consensus provides order fairness or not.
The only related work on verifiability is presented by Li et al. [10], where they provide a protocol to enable message senders to verify whether their messages are correctly processed and ordered by the nodes using an extra round of communication for nodes to send receipts to message senders and use TEE in the message pool to run ordering function. Still, this proposal is costly in communication and it only prevents messages from being held by the nodes but does not prove order fairness. How to provide verifiable fairness efficiently is yet an important topic to explore.
VI-C Adding Flexibility to Order-Insensitive Messages
@@
The Sui blockchain [77] pictured the scenario that in hybrid blockchains some messages do not need to go through consensus if they are accessing a state controlled by a single party and designed a fast commit path for such messages. Another approach [85] is to introduce a consensus that includes a new type of token that can be paid to prioritize messages, allowing order-sensitive messages to be prioritized at a higher cost.
@@
In fact, even the definition of order-sensitive messages is unclear yet. Amiri et al. [26] define order-sensitive messages as those that require accessing a shared resource, where the resulting state changes if the execution order changes. However, messages that also require accessing shared resources may also be order-insensitive. Here we present a potential idea to distinguish order-sensitive messages: instead of asking the system to decide which messages are order-sensitive, ask message senders to define whether their messages are order-sensitive or not. This reduces the work of the system and results in a more precise separation of order-sensitive messages. However, this is only an initial idea we left as future work to further explore.
VI-D Message Ordering As A Service
@@
When implementing message ordering, most existing proposals integrate them into total ordering consensus protocols to minimize the overhead of exchanging the metadata and verifying the ordering. Concurrent with the raising awareness of fair ordering, there is a trend to isolate, and potentially outsource, the process of message sequencing from reaching consensus in distributed ledger systems [86, 76, 87, 88, 89]. In such a design, there are specialized resources that receive messages from users and sequence the messages, before they are passed to the validating nodes to commit to the network.
@@
This can potentially lead to a new design paradigm for achieving order fairness in distributed ledgers. Distributed fair ordering can be taken as an isolated service that serves multiple distributed ledgers. A related proposal is found in an advancement idea for Ethereum [90], which proposed to enable faster execution of ordering in consensus by separating block proposal and transaction ordering [90]. This is also similar to the concept of modular blockchain [91], which separates a blockchain system into multiple sub-systems. From this perspective, a fair-ordering service and be viewed as a sub-system in a modular blockchain, which can also serve multiple consensus and application layer sub-systems. However, there are no other works as far as we know that explore the design of an outsourced fair ordering system.
VI-E Adapting to Permissionless Setting
Most of the existing proposals (except two of them [52, 22]) are in permissioned network settings. In the context of blockchain systems with the need for decentralization, permissionless setting is a major trend. In a permissionless setting, fair ordering is more difficult because, first, there are usually more nodes in such systems in practice, where messages are usually transmitted by gossiping instead of broadcasting. This increases the costs of local-state synchronization. Second, since nodes can freely enter and exit, older messages may be missing in the view of new nodes and expected views from exited nodes may be missing as the required metadata. Therefore, FIFO order in the permissionless setting remains an open challenge to be explored.
VII Conclusion
@@
This paper explores fair message ordering in distributed ledgers, addressing a critical need for fairness in distributed ledgers where manipulation of transaction order, such as Maximal Extractable Value (MEV), becomes a prominent problem. By examining FIFO, random, and blind ordering, we outline the challenges and trade-offs of implementing fair ordering in Byzantine fault-tolerant settings. Our proposed Fair Consensus Factory framework enables flexible integration of fairness into consensus protocols and was demonstrated through optimization of the FIFO ordering consensus protocol, Themis, to reduce latency. This approach shows the potential of modular fairness frameworks to improve both performance and equity in distributed systems. Having a better understanding of fair message ordering accelerates the adoption of blockchain technologies. Our finding provides good guidance for researchers in this area by providing a systematic understanding of existing works and identifying future research directions.
VIII Acknowledgement

### v3

@@ 4.3 Application to Leader-Based and Leaderless Protocols
Figure 5: Transformation of leader-based (top) and leaderless (bottom) consensus via the Fair Consensus Factory
4.4 Preservation of Safety and Liveness
The Fair Consensus Factory transformation preserves the safety and liveness properties of the underlying consensus protocol.
Safety
Safety in consensus protocols ensures that no two honest nodes commit different outputs. Our transformation maintains the original protocol’s agreement mechanism, such as voting or quorum certificates, which guarantees this property.
Liveness
If a Byzantine node sends inconsistent views of $(S,D)$ to different honest nodes during synchronization, it may cause the nodes to compute divergent ordered outputs. Since the agreement logic remains unmodified, safety is preserved, i.e., no two honest nodes will commit different outputs. However, this divergence can prevent the protocol from reaching consensus in the current round, potentially leading to the failure of agreement for this consensus round.
However, this is not a liveness concern as it is mitigable in the next consensus round. In the next round, honest nodes can reshare local views shared in the last round during state synchronization, and additionally exchange hashes of the local input views $h(S_{i},D_{i})$ they receive from other nodes in the previous round. If a node is found to have equivocated, its input can be excluded from the new round. This restores consistency of the locally derived proposal, and allows the protocol to proceed normally.
4.5 Practical Considerations and Limitations
A key limitation arises when messages are not sufficiently disseminated before consensus. If a message $m$ exists in fewer than $f+1$ honest nodes’ local pools, Byzantine nodes may suppress $m$ during synchronization or lie about its metadata, breaking the fairness guarantees.
To mitigate this, some protocols [1, 21] require that clients send messages to all nodes (Figure 6), or assume bounded message propagation delay [19]. Another partial mitigation is to apply blind ordering: nodes are required to execute $r$ without knowing message content, removing the incentive for MEV-style reordering even if unfair metadata is injected.
Figure 6: Two message dissemination modes: one-to-one (left), and broadcast-to-all (right)
Finally, we note that recent proposals suggest relying on trusted execution environments (TEEs) [54, 73, 74] or zero-knowledge proofs [1] to enforce ordering from a single node without requiring all-to-all synchronization. While promising, these approaches introduce trust or computational costs that may conflict with decentralization or latency goals  [^1] With zero-knowledge proofs, the node responsible for applying the ordering rule is also required to generate proof of the computation, which is computationally expensive and therefore results in high latency. To the best of our knowledge, no experimental evaluations have demonstrated otherwise.. For this reason, our proposed framework prioritizes solutions based on all-to-all synchronization and deterministic execution.
5 Case Study: Themis
@@
Themis [1] is a consensus protocol designed to achieve fair ordering through a FIFO-via-voting mechanism. It has been recognized as a promising solution to the blockchain order-fairness problem [13, 11, 75, 76]. In particular, Chen et al. [76] proposed execution-time optimizations for Themis’s ordering algorithm. In this section, we revisit Themis through the lens of the Fair Consensus Factory, illustrating how the framework enables systematic performance improvements while preserving the fairness guarantees of the original design.
5.1 Design of Themis
Themis implements batch order fairness using a FIFO-via-voting algorithm. In each round, given a collection of messages and metadata about their local arrival orders from $t$ nodes, the algorithm constructs a directed graph that captures precedence constraints between messages. The algorithm then produces an ordered batch of messages consistent with this precedence structure. Importantly, unlike earlier work [19], Themis does not require observing a complete Condorcet cycle to output a valid ordering, thus improving responsiveness.
To embed this ordering process in a consensus protocol, Themis builds on HotStuff [70]. Each round, the leader collects local ordering metadata from all nodes, executes the ordering algorithm, and proposes the resulting message batch. To ensure correctness, the proposal includes all metadata so that followers can recompute and verify the leader’s output before voting.
Figure 7 illustrates this process. During the commit phase of the previous round, each node sends its local ordering state to the next-round leader. The leader then runs the FIFO-via-voting algorithm and proposes an ordered batch, along with the metadata used to compute it. All nodes independently re-execute the algorithm and validate the result before proceeding with the HotStuff voting phases.
Figure 7: Themis round structure: the leader gathers local states in the commit phase of the previous round, runs the ordering algorithm, and proposes the ordered batch. The followers re-execute the algorithm to verify the proposal.
5.2 An Alternative Themis Design with Latency Optimization
With Fair Consensus Factory, we propose an alternative design of Themis that still uses $P_{\text{Hotstuff}}$ and $r_{\text{Themis}}$. We show that this alternative design achieves lower latency than the original Themis protocol without sacrificing throughput.
Figure 8 visualizes the resulting protocol after applying the leader-based protocol transformation presented in Section 4.3. This alternative protocol has two key differences compared to the original Themis protocol. First, Themis uses a two-stage synchronization (all-to-leader, leader-to-all), while the alternative protocol uses a single-stage all-to-all synchronization. Second, in Themis, rule execution is separated from rule verification (as shown in Figure 7); in the alternative protocol, every node executes the ordering algorithm locally in the same round.
Figure 8: Optimized Themis structure: nodes exchange metadata in a single all-to-all stage and locally compute the ordering result.
As discussed in Section 4.4, the optimized Themis protocol preserves the safety and liveness guarantee from Hotstuff.
5.3 Experimental Evaluation
We implemented both the original and optimized protocols using a Rust-based HotStuff implementation [77]. To isolate communication and synchronization effects, the running time of the ordering algorithm is abstracted as a fixed 100 ms delay. We measure throughput (transactions per second) and latency (from the time a transaction is sent to a network node to the time the transaction is committed) across 4, 7, 10, and 20 node configurations in a local network, running each configuration for 10 minutes and repeating it three times to ensure stability.
Figure 9: Throughput and latency comparison between original and optimized Themis protocols.
As shown in Figure 9, throughput remains comparable across both versions. However, the optimized protocol achieves a significant reduction in latency, approximately 40% lower, demonstrating the effectiveness of early rule execution and symmetric communication. Importantly, the result shows that the all-to-all synchronization pattern has little negative impact on latency and throughput even as the number of nodes grows. This result supports the claim that the Fair Consensus Factory enables systematic construction and optimization of fair consensus protocols.
6 Fair Ordering Research Roadmap
@@
6.1 How to Choose Fair Ordering Rules
With these categories of fair ordering rules, there are few discussions on which ordering rule is the “better” one, or which one is more suitable for distributed ledgers serving which applications. A possibility is that the ordering rules should be application-specific [78], e.g., first-come-first-served is more suitable for financial applications [4]. Random ordering approach could be the alternative that is a more specific solution to the problem of MEV with a lower ordering algorithm complexity. However, outside financial applications, it is argued that the current most popular fee-based prioritization approach [79] results in higher overall social welfare [80] compared to FIFO.
6.2 Adding Flexibility to Order-Insensitive Messages
@@
The Sui blockchain [72] pictured the scenario that in hybrid blockchains some messages do not need to go through consensus if the resulting state changes only affect the message senders themselves and designed a fast commit path for such messages. Another approach [81] is to introduce a consensus that includes a new type of token that can be paid to prioritize messages, allowing order-sensitive messages to be prioritized at a higher cost.
Amiri et al. [24] define order-sensitive messages as those that require accessing a shared resource, where the resulting state changes if the execution order changes. However, messages that also require accessing shared resources may also be order-insensitive.
@@
Here we present a potential way to distinguish order-sensitive messages: instead of having the system decide which messages are order-sensitive, message senders should define whether their messages are order-sensitive or not. Given such input from message senders, the consensus protocol can apply fair ordering rules only to the ones that are flagged order-sensitive. This reduces the work of the system and results in a more precise separation of order-sensitive messages. However, this is only an initial idea we left as future work to further explore.
6.3 Message Ordering as a Service
@@
When implementing message ordering, most existing proposals integrate them into total ordering consensus protocols to minimize the overhead of exchanging the metadata and verifying the ordering. In line with the rising awareness of fair ordering, there is a trend to isolate, and potentially outsource, the process of message sequencing from reaching consensus in distributed ledger systems [82, 71, 83, 84, 85]. In such a design, specialized resources are used to receive messages from users and sequence the messages before they are passed to the validating nodes to commit to the network.
@@
This can potentially lead to a new design paradigm for achieving order fairness in distributed ledgers. Distributed fair ordering can be taken as an isolated service that serves multiple distributed ledgers. A related proposal is found in an advancement idea for Ethereum [86], which proposed to enable faster execution of ordering in consensus by separating block proposal and transaction ordering [86]. This is also similar to the concept of modular blockchain [87], which separates a blockchain system into multiple sub-systems. From this perspective, a fair-ordering service can be viewed as a standalone system in a modular blockchain, which can also serve multiple consensus and application layer sub-systems. However, there are no other works, as far as we know, that explore the design of an outsourced fair ordering system.
7 Conclusion
@@
This paper explores fair message ordering in distributed ledgers, addressing a critical need for fairness in distributed ledgers where manipulation of transaction order, such as Maximal Extractable Value, becomes a prominent problem. By examining FIFO, random, and blind ordering, we outline the challenges and trade-offs of implementing fair ordering in Byzantine fault-tolerant settings. Our proposed Fair Consensus Factory framework enables flexible integration of fairness into consensus protocols and was demonstrated through redesigning the FIFO ordering consensus protocol, Themis, to reduce latency. This approach shows the potential of modular fairness frameworks to improve both performance and equity in distributed systems. Having a better understanding of fair message ordering accelerates the adoption of blockchain technologies. Our finding provides good guidance for researchers in this area by providing a systematic understanding of existing works and identifying future research directions.
Acknowledgments